solution containing 133 g of magnesium chloride, which has a molar mass of 95.21 g/mol,is dissolved in 0.5 kg of water. Determine the boiling point  and freezing point of the solution. Kb of solvent is 0.52 ⁰𝐶/𝑚, Kf of solvent is 1.86 ⁰𝐶/𝑚


1
Expert's answer
2021-11-25T01:26:59-0500

133/95.21=1.4moles

1= 1.4×0.5×(x-0.52)

1=0.7×(x-0.52)

1.42=x-0.52

X= 1.94°c